# Towards Scalable Language-Image Pre-training for 3D Medical Imaging

### v2

@@ Abstract
The scalability of current language-image pre-training for 3D medical imaging, such as CT and MRI, is constrained by the need for radiologists to manually curate raw clinical studies. In this work, we pioneer pre-training directly on uncurated studies, which both aligns more closely with the radiologist’s workflow and provides a natural path to scalability. However, the unique structure of such data presents new challenges for existing model architectures, which were originally designed for 2D slices or single 3D scans. To address this, we introduce a novel hierarchical attention mechanism inspired by the intrinsic hierarchy of radiology data: slice, scan, and study. We denote our framework as Hierarchical attention for Language-Image Pre-training (HLIP). Trained on 220K studies with 3.13 million scans for brain MRI and 240K studies with 1.44 million scans for head CT, HLIP achieves state-of-the-art performance, e.g., +10.5% balanced ACC on the proposed publicly available brain MRI benchmark Pub-Brain-5; +8.3% and +1.7% macro AUC on head CT benchmarks CQ500 and RSNA, respectively. HLIP also exhibits strong generalizability on existing 3D medical language-image pre-training benchmarks, e.g., +4.3% macro AUC on the Rad-ChestCT benchmark when pre-trained on CT-RATE. These results demonstrate that, with HLIP, directly pre-training on uncurated clinical datasets is a scalable and effective direction for language-image pre-training in 3D medical imaging. The code is available at https://github.com/Zch0414/hlip
@@ 1 Introduction
Language-supervised pre-training is well-suited for radiology, where each study comprises medical images paired with a corresponding radiologist’s report. This natural alignment between visual and textual information has motivated the adaptation of language-image pre-training methods, such as CLIP [1, 2], to learn clinically meaningful radiology representations. CLIP-based models are especially notable for their clinical utility: they demonstrate strong zero-shot transfer performance on diagnostic tasks [3, 4], and their encoders consistently improve performance on multimodal learning benchmarks [5, 6]. In the domain of 2D medical imaging, such as chest X-rays, language-supervised pre-training is a key driver for integrating computer vision into clinical workflows [7, 8, 9].
However, language-image pre-training in 3D medical imaging has yet to reach the scale or performance demonstrated in 2D modalities. For instance, chest X-ray CLIP models have been trained on a 500K corpus [8, 9, 10], achieving human-level performance on multiple diagnostic tasks [9]. In contrast, progress in 3D medical imaging remains limited, with existing models underperforming relative to their 2D counterparts [3, 5]. We attribute the performance gap between 2D and 3D language-image pre-training to two primary factors: data annotation that constrains the training scale, and architectural limitations arising from the complexity of 3D medical imaging.
Computed tomography (CT) and magnetic resonance imaging (MRI) generate 3D volumetric images across various anatomical regions, including the brain, chest, and abdomen. As illustrated in Figure 1, a standard MRI study typically includes several sequences (e.g., T1-weighted, T2-weighted, and FLAIR), each contributing distinct diagnostic information. Similarly, CT studies often include scans acquired with varying orientations or scanner settings within the same study. To perform language-image pre-training for such data, a common strategy is to curate datasets by having radiologists manually select a representative scan or slice from each study, as shown in Figure 1(a) [4, 3, 5, 6], which presents a significant barrier to the scalability. In contrast, pre-training on uncurated studies aligns more closely with real-world practice and readily expands the data scale, as it imposes no additional burden on radiologists. Despite this, the unique structure of such data presents new challenges for current visual encoders, which were originally designed for 2D images or single 3D scans [11, 12, 13]. As shown in Figure 1(b), even trained on a large-scale domain-specific dataset, naively encoding uncurated studies with the Vision Transformer (ViT) [11] results in only comparable zero-shot transferability to the state-of-the-art (SOTA) biomedical CLIP [14] trained on the PubMed corpus. Particularly, encoding the entire study can produce tokens on the order of $10^{4}$, which both incurs substantial computational overhead and limits performance [15].
@@
In this work, we first address the key barriers to scaling language-image pre-training for 3D medical imaging by pioneering the use of uncurated studies. Second, to effectively extract features from such data, we introduce a novel hierarchical attention mechanism leverages the natural hierarchy of radiology data: slice, scan, and study. We name this framework Hierarchical attention for Language-Image Pre-training (HLIP). Unlike architectural designs such as Swin [12], MViT [16, 17], and Hiera [13], HLIP leverages the inherent data structure to define the attention scope, making it suitable for uncurated studies that contain multiple 3D scans. Compared to window attention that captures only local features, slice or scan attention can capture all diagnostic features while also providing constructive priors for learning study representations. Moreover, slice and scan attention is already much lighter than study attention, and such minimal adaptation of the original ViT remains orthogonal to flash attention [18] and patch dropout [19], further reducing the computational burden.
@@
Trained on our health system, HLIP outperforms the SOTA [14] on the proposed publicly available brain MRI benchmark, Pub-Brain-5, by 10.5% balanced ACC; and surpasses the head CT foundation model [20] by 8.3% and 1.7% macro AUC on the CQ500 [21] and RSNA [22] benchmarks, respectively. HLIP also demonstrates strong generalizability on the curated 3D medical language-image pre-training benchmark CT-RATE [3], which contains only one scan per study, outperforming the SOTA [23] by 4.3% macro AUC on the external evaluation Rad-ChestCT [24].
Our paper makes the following contributions:
@@ 2 Related Work
Language-Image Pre-training in Radiology has been developed for 2D and 3D medical imaging, facilitated by public datasets such as CheXpert [25], MIMIC-CXR [26], and CT-RATE [3]. In 2D imaging, techniques such as local alignment [27, 28, 29], knowledge enhancement [8, 30], and longitudinal analysis [31] have been explored. CheXZero [9] demonstrates strong empirical results comparable to those of human experts in the domain of chest X-rays. BiomedCLIP [32] and ConceptCLIP [14] have been trained on more than 15 million sample pairs, achieving strong performance across modalities, including radiology. In addition, other works [7, 10, 33] have also contributed significant insights and achieved impressive performance. 3D medical imaging offers a more comprehensive view of anatomical structures. However, due to its computational cost and limited data availability, several studies [34, 35, 36] focus on bridging the domain gap between 2D and 3D. For example, UniMedI [36] learns a shared feature space for both 2D and 3D modalities and demonstrates improvements across both. BIUD [34] distills 3D representations from a well-trained 2D model [9], significantly improving data efficiency.
@@
More recently, aided by advances in hardware and the availability of public datasets [3], several studies [3, 5, 6, 37] have performed language-image pre-training on 3D imaging, surpassing methods [34] that still rely on 2D representations. Specifically, CT-CLIP [3] has been pre-trained on 20,000 paired chest CT scans and reports. M3D [5] explores a versatile multi-modal large language model designed for universal 3D medical imaging analysis. fVLM [6] proposes a fine-grained pre-training framework that relies on segmentation to perform organ alignment. However, while these methods have demonstrated promising performance on various tasks such as zero-shot abnormality detection and report generation, several factors limit their training scalability and real-world applicability. For example, CT-CLIP and M3D are constrained to curated studies containing only a single imaging scan. Such studies rely heavily on human annotation and also fail to reflect real-world scenarios, where studies typically include multiple scans. fVLM further depends on a segmentation model, which introduces bias from segmentation quality and largely limits its scalability and generalizability. Moreover, these methods suffer from inefficient modeling of 3D medical imaging, leading to small batch sizes (e.g., 48), which are insufficient for effective language-image pre-training [2].
@@
Efficient Language-Image Pre-training is crucial, as it directly affects the number of sample pairs seen during training and contrasted per batch, two key factors of model capacity. Existing hierarchical architectures [12, 3] have been widely adopted to improve efficiency. However, as analyzed in Appendix D.2, these models may be less efficient than commonly assumed in the context of 3D medical imaging, even when compared with the original ViT. From the architecture perspective, components such as relative position embedding are expensive for 3D inputs and is not compatible with recent advancements like flash attention [18]. From the training perspective, FLIP [19] demonstrates a favorable trade-off by randomly removing 50% of tokens during training. However, models that require a fixed activation shape [12, 17, 38, 13] cannot benefit from this strategy. Moreover, many works in radiology [39, 9, 40, 6] have demonstrated the benefits of the universal feature learned by MAE [41]. Therefore, modeling 3D medical imaging with minimal adaptation of the original ViT [11] is a promising research direction and stands to benefit from recent advancements such as flash attention, patch dropout, and pre-trained models.
@@ 3 Method
Our goal is to perform language-image pre-training on uncurated studies. Given a study $S\in\mathbb{R}^{M\times 1\times D\times H\times W}$, where $M$ denotes the number of single-channel 3D scans, and $D$, $H$, and $W$ represent the depth, height, and width dimensions of each scan, respectively. Following the ViT [11], each scan is divided into a grid of non-overlapping volumes of size $(\frac{D}{d},\frac{H}{h},\frac{W}{w})$. These volumes are projected into visual tokens $F_{s}\in\mathbb{R}^{N\times c}$, where $N=M\times d\times h\times w$ is the total number of tokens and $c$ denotes the number of channels. As $N$ can be on the order of $10^{4}$, computing self-attention over all tokens throughout the ViT backbone is prohibitive in memory and also limits performance [15]. We introduce a novel hierarchical attention mechanism guided by the inherent data hierarchy of uncurated studies, enabling lighter self-attention while introducing effective priors derived from this hierarchy: slice, scan, and study.
@@ 3.1 Hierarchical Attention Mechanism
  • Study attention computes a single self-attention operation over all $N$ tokens in a study. The I/O complexity [42] of the study attention is $\Omega(N^{2}+N\times c)$ [18], as detailed in Appendix D.
@@
The proposed hierarchical attention mechanism contrasts with existing methods that modify attention computation using convolution [16, 17], pooling [13], or shifted windows [12], all of which require either regular activation shapes or attention masks. Our mechanism relies solely on a simple reshape operation, as there is no overlap between different scans or groups of adjacent slices [^1] With the original attention layer unchanged, study attention takes the input of size $(B,\ M\times d\times h\times w,\ c)$; scan attention takes the input of size $(B\times M,\ d\times h\times w,\ c)$; and slice attention takes the input of size $(B\times M\times d,\ h\times w,\ c)$, where $B$ denotes the batch size., allowing greater flexibility for uncurated studies that comprise multiple scans. As shown in Figure 2, each layer in the ViT can perform attention at any level. In practice, we evenly divide the ViT backbone into four subsets of layers (e.g., three layers per subset for the 12-layer ViT-B) and apply study attention only to the last layer of each subset, while the remaining layers perform the lighter scan or slice attention. Moreover, this intentionally simple design also enables seamless integration with recent efficiency advances such as patch dropout [19] and flash attention [18].
@@ 3.2 Implementation
Model. Our visual encoder is a MAE pre-trained ViT-B [41]. The input is rescaled to [0,1] and normalized with MAE’s averaged mean and std. Compared to the original ViT designed for 2D RGB images, our vision encoder only has three differences: (1) each scan is divided into 3D volumes instead of 2D patches; (2) the positional embedding encodes 3D spatial coordinates within each scan, along with an additional dimension to distinguish different scans; The cls token is minimally adapted to propagate information across layers that perform different hierarchies of attention.
@@
For uncurated datasets, we use scan attention with 4 evenly distributed study attentions. A curated dataset with only one scan per study corresponds to the special case of M=1 in our scenario. For such datasets, we employ slice attention with 4 evenly distributed scan attentions; We use PubMedBERT [44] as our text encoder. Model configurations are provided in Appendix B.
@@
Preprocessing. We do not standardize the orientation or spacing. Instead, we consistently align the depth dimension with the through-plane axis of each scan, and then resize it to a fixed shape of (48,224,224). As this differs from prior practice [4, 3, 6], we provide a discussion in Appendix A.3. For brain MRI, we apply [0.5,99.5] percentile clipping to the intensity values. For head CT, we expand each scan into three separate scans, with Hounsfield Unit (HU) values truncated to [0,120] for soft tissue, [-20,180] for blood vessels, and [-800,2000] for bone.
Pre-training. Our implementation builds upon OpenCLIP [45] and FLIP [19]. We apply 25% patch dropout by default as regularization and acceleration. We do not apply additional augmentation or unmasked fine-tuning. Training 20 epochs on our uncurated datasets takes ˜1 day with a batch size of 256 on 8 L40 GPUs. More details are provided in Appendix B.2.

### v3

@@ Abstract
The scalability of current language-image pre-training for 3D medical imaging, such as CT and MRI, is constrained by the need for radiologists to manually curate raw clinical studies. In this work, we pioneer pre-training directly on uncurated studies, which both aligns more closely with the clinical workflow and provides a natural path to scalability. However, the unique structure of such data presents new challenges for existing model architectures, which were originally designed for 2D slices or single 3D scans. To address this, we introduce a novel hierarchical attention mechanism inspired by the intrinsic hierarchy of radiology data: slice, scan, and study. We denote our framework as Hierarchical attention for Language-Image Pre-training (HLIP). Trained on 220K studies with 3.13 million scans for brain MRI and 240K studies with 1.44 million scans for head CT, HLIP achieves state-of-the-art performance, e.g., +10.5% balanced ACC on the proposed publicly available brain MRI benchmark Pub-Brain-5; +8.3% and +1.7% macro AUC on head CT benchmarks CQ500 and RSNA, respectively. HLIP also exhibits strong generalizability on existing 3D medical language-image pre-training benchmarks, e.g., +4.3% macro AUC on the Rad-ChestCT benchmark when pre-trained on CT-RATE. These results demonstrate that, with HLIP, directly pre-training on uncurated clinical datasets is a scalable and effective direction for language-image pre-training in 3D medical imaging. The code is available at https://github.com/zch0414/hlip
@@ 1 Introduction
Language-supervised pre-training is well-suited for radiology, where each study comprises medical images paired with a corresponding radiologist’s report. This natural alignment between visual and textual information has motivated the adaptation of language-image pre-training methods, such as CLIP (Zhang et al., 2022a; Radford et al., 2021), to learn clinically meaningful radiology representations. CLIP-based models are especially notable for their clinical utility: they demonstrate strong zero-shot transfer performance on diagnostic tasks (Hamamci et al., 2024; Blankemeier et al., 2024), and their encoders consistently improve performance on multimodal learning benchmarks (Bai et al., 2024; Shui et al., 2025). In the domain of 2D medical imaging, such as chest X-rays, language-supervised pre-training is a key driver for integrating computer vision into clinical workflows (Boecking et al., 2022; Wang et al., 2022b; Tiu et al., 2022).
However, language-image pre-training in 3D medical imaging has yet to reach the scale or performance demonstrated in 2D modalities. For instance, chest X-ray CLIP models have been trained on a 500K corpus (Wang et al., 2022b; Tiu et al., 2022; You et al., 2023), achieving human-level performance on multiple diagnostic tasks (Tiu et al., 2022). In contrast, progress in 3D medical imaging remains limited, with existing models underperforming relative to their 2D counterparts (Hamamci et al., 2024; Bai et al., 2024). We attribute the performance this gap to two primary factors: data annotation that constrains the training scale, and architectural limitations arising from the complexity of 3D medical imaging.
Computed tomography (CT) and magnetic resonance imaging (MRI) generate 3D volumetric images across various anatomical regions, including the brain, chest, and abdomen. As illustrated in Figure 1, a standard MRI study typically includes several sequences (e.g., T1-weighted, T2-weighted, and FLAIR), each contributing distinct diagnostic information. Similarly, CT studies often include scans acquired with varying orientations or scanner settings within the same study. To perform language-image pre-training for such data, a common strategy is to curate datasets by having radiologists manually select a representative scan or slice from each study, as shown in Figure 1(a) (Blankemeier et al., 2024; Hamamci et al., 2024; Bai et al., 2024; Shui et al., 2025), which presents a significant barrier to the scalability. In contrast, pre-training on uncurated studies aligns more closely with real-world practice and readily expands the data scale, as it imposes no additional burden on radiologists. Despite this, the unique structure of such data presents new challenges for current visual encoders, which were originally designed for 2D images or single 3D scans (Dosovitskiy et al., 2020; Liu et al., 2021; Ryali et al., 2023). As shown in Figure 1(b), even trained on a large-scale domain-specific dataset, naively encoding uncurated studies with the Vision Transformer (ViT) (Dosovitskiy et al., 2020) results in only comparable zero-shot transferability to the state-of-the-art (SOTA) biomedical CLIP (Nie et al., 2025) trained on the PubMed corpus. Particularly, encoding the entire study can produce tokens on the order of $10^{4}$, which both incurs substantial computational overhead and limits performance (Barbero et al., 2024).
@@
In this work, we first address the key barriers to scaling language-image pre-training for 3D medical imaging by pioneering the use of uncurated studies. Second, to effectively extract features from such data, we introduce a novel hierarchical attention mechanism leverages the natural hierarchy of radiology data: slice, scan, and study. We name this framework Hierarchical attention for Language-Image Pre-training (HLIP). Unlike architectural designs such as Swin (Liu et al., 2021), MViT (Fan et al., 2021; Li et al., 2022b), and Hiera (Ryali et al., 2023), HLIP leverages the inherent data structure to define the attention scope, making it suitable for uncurated studies that contain multiple 3D scans. Compared to window attention that captures only local features, slice or scan attention can capture all diagnostic features while also providing constructive priors for learning study representations. Moreover, slice and scan attention is already much lighter than study attention, and such minimal adaptation of the original ViT remains orthogonal to flash attention (Dao et al., 2022) and patch dropout (Li et al., 2023), further reducing the computational burden.
@@
Trained on our health system, HLIP outperforms the SOTA (Nie et al., 2025) on the proposed publicly available brain MRI benchmark, Pub-Brain-5, by 10.5% balanced ACC; and surpasses the head CT foundation model (Zhu et al., 2025) by 8.3% and 1.7% macro AUC on the CQ500 (Chilamkurthy et al., 2018) and RSNA (Flanders et al., 2020) benchmarks, respectively. HLIP also demonstrates strong generalizability on the curated 3D medical language-image pre-training benchmark CT-RATE (Hamamci et al., 2024), which contains only one scan per study, outperforming the SOTA (Liu et al., 2023) by 4.3% macro AUC on the external evaluation Rad-ChestCT (Draelos et al., 2021). Our paper makes the following contributions:
@@ 2.1 Language-Image Pre-training in Radiology
has been developed for 2D and 3D medical imaging, facilitated by public datasets such as CheXpert (Irvin et al., 2019), MIMIC-CXR (Johnson et al., 2019), and CT-RATE (Hamamci et al., 2024). In 2D imaging, techniques such as local alignment (Huang et al., 2021; Wang et al., 2022a; Müller et al., 2022), knowledge enhancement (Wang et al., 2022b; Wu et al., 2023), and longitudinal analysis (Bannur et al., 2023) have been explored. CheXZero (Tiu et al., 2022) demonstrates strong empirical results comparable to those of human experts in the domain of chest X-rays. BiomedCLIP (Zhang et al., 2023a) and ConceptCLIP (Nie et al., 2025) have been trained on more than 15 million sample pairs, achieving strong performance across modalities, including radiology. In addition, other works (Boecking et al., 2022; You et al., 2023; Zhang et al., 2023b) have also contributed significant insights and achieved impressive performance.
3D medical imaging offers a more comprehensive view of anatomical structures. However, due to its computational cost and limited data availability, several studies (Cao et al., 2024; Wang et al., 2024; He et al., 2024) focus on bridging the domain gap between 2D and 3D. For example, UniMedI (He et al., 2024) learns a shared feature space for both 2D and 3D modalities and demonstrates improvements across both. BIUD (Cao et al., 2024) distills 3D representations from a well-trained 2D model (Tiu et al., 2022), significantly improving data efficiency.
@@
More recently, aided by advances in hardware and the availability of public datasets (Hamamci et al., 2024), several studies (Hamamci et al., 2024; Bai et al., 2024; Shui et al., 2025; Lai et al., 2025) have performed language-image pre-training on 3D imaging, surpassing methods (Cao et al., 2024) that still rely on 2D representations. Specifically, CT-CLIP (Hamamci et al., 2024) has been pre-trained on 20,000 paired chest CT scans and reports. M3D (Bai et al., 2024) explores a versatile multi-modal large language model designed for universal 3D medical imaging analysis. fVLM (Shui et al., 2025) proposes a fine-grained pre-training framework that relies on segmentation to perform organ alignment. However, while these methods have demonstrated promising performance on various tasks such as zero-shot abnormality detection and report generation, several factors limit their training scalability and real-world applicability. For example, CT-CLIP and M3D are constrained to curated studies containing only a single imaging scan. Such studies rely heavily on human annotation and also fail to reflect real-world scenarios, where studies typically include multiple scans. fVLM further depends on a segmentation model, which introduces bias from segmentation quality and largely limits its scalability and generalizability. Moreover, these methods suffer from inefficient modeling of 3D medical imaging, leading to small batch sizes (e.g., 48), which are insufficient for effective language-image pre-training (Radford et al., 2021).
2.2 Efficient Language-Image Pre-training
@@
is crucial, as it directly affects the number of sample pairs seen during training and contrasted per batch, two key factors of model capacity. Existing hierarchical architectures (Liu et al., 2021; Hamamci et al., 2024) have been widely adopted to improve efficiency. However, as analyzed in Appendix E.2, these models may be less efficient than commonly assumed in the context of 3D medical imaging, even when compared with the original ViT. From the architecture perspective, components such as relative position embedding are expensive for 3D inputs and is not compatible with recent advancements like flash attention (Dao et al., 2022). From the training perspective, FLIP (Li et al., 2023) demonstrates a favorable trade-off by randomly removing 50% of tokens during training. However, models that require a fixed activation shape (Liu et al., 2021; Li et al., 2022b; a; Ryali et al., 2023) cannot benefit from this strategy. Moreover, many works in radiology (Matsoukas et al., 2022; Tiu et al., 2022; Zhao and Shen, 2024; Shui et al., 2025) have demonstrated the benefits of the universal feature learned by MAE (He et al., 2022). Therefore, modeling 3D medical imaging with minimal adaptation of the original ViT (Dosovitskiy et al., 2020) is a promising research direction and stands to benefit from recent advancements such as flash attention, patch dropout, and pre-trained models.
@@ 3 Method
Our goal is to perform language-image pre-training on uncurated studies. Given a study $S\in\mathbb{R}^{M\times 1\times D\times H\times W}$, where $M$ denotes the number of single-channel 3D scans, and $D$, $H$, and $W$ represent the depth, height, and width dimensions of each scan, respectively. Following the ViT (Dosovitskiy et al., 2020), each scan is divided into a grid of non-overlapping volumes of size $(\frac{D}{d},\frac{H}{h},\frac{W}{w})$. These volumes are projected into visual tokens $F_{s}\in\mathbb{R}^{N\times c}$, where $N=M\times d\times h\times w$ is the total number of tokens and $c$ denotes the number of channels. As $N$ can be on the order of $10^{4}$, computing self-attention over all tokens throughout the ViT backbone is prohibitive in memory and also limits performance (Barbero et al., 2024). We introduce a novel hierarchical attention mechanism guided by the inherent data hierarchy of uncurated studies, enabling lighter self-attention while introducing effective priors derived from this hierarchy: slice, scan, and study.
@@ 3.1 Hierarchical Attention Mechanism
  • Study attention computes a single self-attention operation over all $N$ tokens in a study. The I/O complexity (Aggarwal and Vitter, 1988) of the study attention is $\Omega(N^{2}+N\times c)$ (Dao et al., 2022), as detailed in Appendix E.
@@
The proposed hierarchical attention mechanism contrasts with existing methods that modify attention computation using convolution (Fan et al., 2021; Li et al., 2022b), pooling (Ryali et al., 2023), or shifted windows (Liu et al., 2021), all of which require either regular activation shapes or attention masks. Our mechanism relies solely on a simple reshape operation, as there is no overlap between different scans or groups of adjacent slices [^1] With the original attention layer unchanged, study attention takes the input of size $(B,\ M\times d\times h\times w,\ c)$; scan attention takes the input of size $(B\times M,\ d\times h\times w,\ c)$; and slice attention takes the input of size $(B\times M\times d,\ h\times w,\ c)$, where $B$ denotes the batch size., allowing greater flexibility for uncurated studies that comprise multiple scans. As shown in Figure 2, each layer in the ViT can perform attention at any level. In practice, we evenly divide the ViT backbone into four subsets of layers (e.g., three layers per subset for the 12-layer ViT-B) and apply study attention only to the last layer of each subset, while the remaining layers perform the lighter scan or slice attention. Moreover, this intentionally simple design also enables seamless integration with recent efficiency advances such as patch dropout (Li et al., 2023) and flash attention (Dao et al., 2022).
@@ 3.2 Implementation
Model. Our visual encoder is a MAE pre-trained ViT-B (He et al., 2022). The input is rescaled to [0,1] and normalized with MAE’s averaged mean and std. Compared to the original ViT designed for 2D RGB images, our vision encoder only has three differences: (1) each scan is divided into 3D volumes instead of 2D patches; (2) the positional embedding encodes 3D spatial coordinates within each scan, along with an additional dimension to distinguish different scans; The cls token is minimally adapted to propagate information across layers that perform different hierarchies of attention.
@@
For uncurated datasets, we use scan attention with 4 evenly distributed study attentions. A curated dataset with only one scan per study corresponds to the special case of M=1 in our scenario. For such datasets, we employ slice attention with 4 evenly distributed scan attentions; We use PubMedBERT (Gu et al., 2020) as our text encoder. Model configurations are provided in Appendix B.
@@
Preprocessing. We do not standardize the orientation or spacing. Instead, we consistently align the depth dimension with the through-plane axis of each scan, and then resize it to a fixed shape of (48,224,224). As this differs from prior practice (Blankemeier et al., 2024; Hamamci et al., 2024; Shui et al., 2025), we provide a discussion in Appendix A.3. For brain MRI, we apply [0.5,99.5] percentile clipping to the intensity values. For head CT, we expand each scan into three separate scans, with Hounsfield Unit (HU) values truncated to [0,120] for soft tissue, [-20,180] for blood vessels, and [-800,2000] for bone.
Pre-training. Our implementation builds upon OpenCLIP (Cherti et al., 2023) and FLIP (Li et al., 2023). We apply 25% patch dropout by default as regularization and acceleration. We do not apply additional augmentation or unmasked fine-tuning. Training 20 epochs on our uncurated datasets takes ˜1 day with a batch size of 256 on 8 L40 GPUs. More details are provided in Appendix B.2.